OpenStudy (anonymous):

Find the derivative of f(x)=-(9)/(x) at x=6

Parth (parthkohli):

\[f(x) = \dfrac{-9}{x}\]Is that the function?

OpenStudy (anonymous):

\[f(x)= - \frac{ 9 }{ x }\]

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Yeah, the same thing. Heard of the product rule or the quotient rule? Either will work here.

OpenStudy (anonymous):

Wait parthy..

Parth (parthkohli):

Sure, waiting.

Parth (parthkohli):

Oh yeah!

OpenStudy (anonymous):

\[\frac{d}{dx}(x^n) = n \cdot x^{n-1}\]

Parth (parthkohli):

Yeah, that.

OpenStudy (anonymous):

And remember: \[\frac{1}{x^a} = x^{-a}\]

Parth (parthkohli):

\[f(x) = -9x^{-1}\]We have a winner here.

Parth (parthkohli):

Grr, why did I even mention the Product Rule.

OpenStudy (anonymous):

@ParthKohli @waterineyes I'm lost

Parth (parthkohli):

Have you heard of the Power Rule?\[\dfrac{d}{dx}(x^n) = n x^{n - 1}\]

OpenStudy (anonymous):

@ParthKohli no

OpenStudy (anonymous):

Net got disconnected.. Yeah, we have @ParthKohli here..

Parth (parthkohli):

@Ambition That's sad... :-|

Parth (parthkohli):

What all do you know about differentiation and derivatives?

OpenStudy (anonymous):

@ParthKohli nothing

OpenStudy (anonymous):

@Ambition Sorry I have to ask this question. Why didn't your teacher teach you the power rule?

Parth (parthkohli):

@Ambition I'm afraid I can't help you then. Take a look at your course notes and lectures.

OpenStudy (anonymous):

@BABYShawol184 he was barely in school @ParthKohli thanks for your help

OpenStudy (anonymous):

Wow. Well good luck @ParthKohli helped a lot so you should be good :) Good Luck ^.^

OpenStudy (anonymous):

the derivative of the reciprocal function \[f(x)=\frac{1}{x}\] is\[f'(x)=-\frac{1}{x^2}\] it is a very common function, best to simply memorize it